OpenStudy (livyloo52):

What property is 8+(y+4)=(8+y)+4??

OpenStudy (hope210):

Are there any options?

OpenStudy (livyloo52):

Yes, a associative property, additive identity property, additive inverse property

OpenStudy (yanasidlinskiy):

https://www.quia.com/jg/332021list.html Which one do you think it is? ^^Use the website above I have provided you with.

OpenStudy (kg1975):

m^4 -8m = 0 m(m ³ -8) =0 m( m-2) (m² +2m +4) =0 the quadratic gives the complex roots